# Subsystem Thermalization Hypothesis in Quantum Spin Chains with Conserved Charges

### v1

@@ II Thermalization of Quantum Systems with Conserved Charges
Figure 2: The evolving pattern of one-site entanglement entropy $S[\rho_{1}]$ of a $10$-site nonintegrable XXZ chain specified by (IV) and (22). This model has a $U(1)$ conserved charge denoted by $Q_{1}$, see (23). Each solid line represents the average of $S[\rho_{1}]$ over 500 states at each time step, which are evolved from the initial product states with fixed $\langle Q_{1}\rangle$. In this figure, we consider $\langle Q_{1}\rangle=0.0$ (red), $1.0$ (blue) and $2.0$ (green). The dashed lines are the values of $S[\rho_{1}]$ of the corresponding GGE states. The result shows that values of $S[\rho_{1}]$ saturate to a non-thermal value around $T=25$.
@@
There is no unique definition of a typical state. In this work, we adopt the following definition used in [31]: its entanglement entropy of a single site with the rest of the spin chain, denoted by $S[\rho_{1}]$, is about the maximal value given by the time evolution of an initial product state. This will exclude some transient states, which may not be typical regarding entanglement spreading. We give an example for this construction by considering a ten-site nonintegrable XXZ spin chain with a $U(1)$ conserved charge $Q_{1}$. We then use ED to evolve a randomly generated product state with zero $S[\rho_{1}]$ but fixed $\langle Q_{1}\rangle$. In Fig. 2, we show the evolving pattern of the averaged $S[\rho_{1}]$ over $500$ initial states. The result shows that the (averaged) $S[\rho_{1}]$ values saturate to a maximal after $T\simeq 25$. However, these maximal values may differ from those given by the corresponding GGE states. In this work, the typical states we collect for the demographic of subsystem thermalization hypothesis are obtained by evolving the randomly generated product states with time duration $T=1000$. This should be long enough to ensure the typicality. Based on this, we will prepare the typical states by choosing $T=1000$.
@@
III Thermalization demographics for quantum Ising spin chains
@@
As a starter, we consider the quantum Ising spin chain in the transverse field $h_{x}$ and longitudinal field $h_{z}$, given by the Hamiltonian
$H=\sum_{i=1}^{L}(h_{x}\sigma_{i}^{x}+h_{z}\sigma_{i}^{z}+\sigma_{i}^{z}\sigma_{i+1}^{z}),$ (17)
@@
where $\sigma_{i}^{x}$, $\sigma_{i}^{z}$ are Pauli matrix operator on the site $i$, and $L$ is the number of spins. This model possesses a $Z_{2}$ reflection symmetry $\Pi$, which swaps the order of spins by sending site $i$ to $L+1-i$, i.e.,
@@
$\Pi:=\begin{cases}P_{1,L}P_{2,L-2}\cdots P_{\frac{L}{2},\frac{L+2}{2}}&\quad\text{if }L=\text{ is even}\\ P_{1,L}P_{2,L-2}\cdots P_{\frac{L-1}{2},\frac{L+3}{2}}&\quad\text{if }L=\text{is odd}\end{cases}\qquad$ (18)
where $P_{k,l}=(\sigma_{k}^{x}\sigma_{l}^{x}+\sigma_{k}^{y}\sigma_{l}^{y}+\sigma_{k}^{z}\sigma_{l}^{z}+I)/2$ is the permutation operator and $I$ is the identity operator. $P_{k,l}$ can swap the state of $k^{th}$ and $l^{th}$ sites.
@@
Figure 3: Patterns of typical-state thermalization demographics of $S^{t}_{A=1}$ of a $10$-site Ising chain of (17) near an integrable point when comparing with the corresponding GGE state specified by the chemical potentials $(\beta_{0},\beta_{1})$ of the conserved charges (20). The phase diagram of the Ising chain is shown in the inset of the bottom subfigure, on which a redline is a region we plot for thermalization demographics. The integrable point on the redline is $h_{x}=1.05$ and $h_{z}=0$. We plot $10$ demographics from $h_{z}=0$ to $h_{z}=0.1$. At the integrable point, the GGE state is just a p-GGE one, as all the conserved charges other than $Q_{1}$ are ignored. The results show that the thermalization hypothesis breaks down badly near the integrable point but gradually holds better as moving away from the integrable point.
@@
This model is integrable if $h_{x}=0$ or $h_{z}=0$. Otherwise, it is nonintegrable. Its phase diagram is shown in the inset of the bottom subfigure of Fig. 3, e.g., see [35]. It was shown that the level-spacing statistics of the superselection sectors with $\Pi=\pm 1$ match with the Wigner-Dyson distribution for the nonintegrable Ising chain [31]. Despite that, we can examine how the pattern of the demographics of $S^{t}_{A}$ changes by tuning the coupling constant to make the model more nonintegrable. The thermalization hypothesis holds well if most of the population lives near $S_{A}^{t}=0$. Otherwise, we can quantify the violation of the thermalization hypothesis by the spreading of the population away from $S_{A}^{t}=0$. A typical result for such pattern changes is shown in Fig. 3. We see that the demographics of $S_{1}^{t}$ for the integrable case violate the subsystem thermalization hypothesis badly. However, as the model moves away from the integrable point, the values of $S_{1}^{t}$ gradually converge to zero.
From now on, we will focus on the nonintegrable Ising chains, specifically for the model with the following parameters:
@@
We will present $S_{A}^{t,a,e}$ demographics for various kinds of p-GGE states.
@@ III.1 Demographics of typical-state thermalization
Figure 5: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{0}$ for the Ising chain of (19) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, i.e., $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{1}\rangle$.
@@
We first consider the subsystem thermalization hypothesis for the typical states. In Fig. 4, we show the thermalization demographics of the relative entropy $S^{t}_{A}$ of the region $A$ with $1$, $3$, $5$ sites (denoted by $A=1,3,5$) between a typical state and the corresponding GGE specified by $(\beta_{0},\beta_{1})$. Since the GGE contains two chemical potential parameters $\beta_{0}$ and $\beta_{1}$, thus we present the demographics in the form of density plots. For $A=1$, $S^{t}_{A}$’s of all the typical states considered are almost vanishing. This implies that the thermalization hypothesis for the typical states works well for $A=1$ when compared with the corresponding GGE thermal states. On the other hand, for $A=3$, the thermalization hypothesis still looks good, with some exceptions for large $|\beta_{0}|$ or $|\beta_{1}|$. However, when $A=5$, the thermalization hypothesis breaks down. The above patterns of dependence of validity of the thermalization hypothesis on the size $A$ are expected and also give the explicitly quantitative picture.
@@
Besides checking the typical-state thermalization hypothesis, Fig. 4 also reveals the demographics of the typical states on the $(\beta_{0},\beta_{1})$ of the corresponding GGE ensemble states. Interestingly, the inverse temperature $\beta_{0}$ can be either positive or negative and almost symmetric about $\beta_{0}=0$. The GGE states with negative (inverse) temperatures are hypothetical and cannot be realized naturally. However, we see that they can be realized by typical states. Moreover, the thermalization hypothesis works even for such hypothetical GGE thermal ensemble states. On the other hand, the chemical potential $\beta_{1}$ for reflection symmetry $\Pi$ is almost negative for the typical states we construct. It is unclear why there is such a bias.
@@
After considering the comparison with GGE states, we now compare with the corresponding p-GGE thermal states. In Fig.5, we present the typical-state demographics of $S^{t}_{A}$ when compared with the corresponding p-GGE states specified only by $\beta_{0}$ (inverse temperature). On the other hand, in Fig. 6, we compare the typical states with the corresponding p-GGE states specified only by $\beta_{1}$. We see that the features of demographics of p-GGE on either $\beta_{0}$, i.e., symmetric about $\beta_{0}$, and $\beta_{1}$, i.e., favors negative values, are the same as the ones of GGE. Regarding the validity of the typical-state thermalization hypothesis, we see in both Fig. 5 and 6 the similar patterns of dependence on the size $A$ as in Fig. 4 for GGE states’s comparison. In all three cases, the thermalization hypothesis for the typical states works quite well for $A=1$ and gradually worsens with increasing $A$.
@@
Although it is generally believed that the typical-state thermalization hypothesis will hold for small subsystems, we shall emphasize that it is unclear if it will hold for various thermal ensemble states and how local the subsystem should be for it to hold. Our results presented here have clarified these issues.
@@ III.2 Demographics of almost superselection-state thermalization
In Fig. 7, we consider the almost superselection typical states with fixed $\langle Q_{1}\rangle$ and compare them with the corresponding p-GGE states specified by $\beta_{0}$. In Fig. 8, we swap the role of the $Q_{0}$ and $Q_{1}$, that is, compare the almost superselection states with fixed $\langle Q_{0}\rangle$ to the corresponding p-GGE states specified by $\beta_{1}$. In both cases, we show the population ratio demographics against $S_{A}^{a}$. The population ratio for a given $\langle Q_{a}\rangle$ is defined with respect to the total number of all typical states considered. We can quantify the violation of the thermalization hypothesis by the spreading of the population away from $S_{A}^{a}=0$. In the first case, as shown in Fig. 7, the thermalization hypothesis holds well for each almost superselection sector if $A=1$. Moreover, the patterns of the populations for different $A$’s do not change much as $\langle Q_{1}\rangle$ varies. In the second case, as shown in Fig. 8, the thermalization holds well only for $\langle Q_{0}\rangle=0$ and worsens more and more as $\langle Q_{0}\rangle=0$ increases even for $A=1$. In this case, even the typical-state thermalization holds, but its fine-grained version does not. Unlike in Fig. 7, the patterns of populations for different $A$’s change more significantly as $\langle Q_{0}\rangle$ varies.
It may look strange why the coarse-grained version of the thermalization hypothesis holds, but most of its fined-grained versions fail. However, if we look into the scales of the population ratio in each subfigure, it can be understood that the population respecting the thermalization hypothesis dominates; see also Fig. 6(b), where the sector of $\langle Q_{0}\rangle=0$, which obeys the thermalization hypothesis as shown in the first subfigure of Fig. 8, has the largest population.
@@ III.3 Demographics of eigenstate thermalization
Figure 9: Subsystem eigenstate thermalization demographics: $S^{e}_{A=1,3,5}$ v.s. $\beta_{0}$ for the Ising chain of (19) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, i.e., $A=3$ (blue circles) and $A=5$ (black triangles).
Figure 10: Subsystem eigenstate thermalization demographics of $S^{e}_{A=1}$ for the Ising chain of (19) by comparing with (i) GGE states (empty red) and (ii) p-GGE states specified by $\beta_{1}$ (solid pink). We also present the typical-state counterparts (empty blue and solid gray) for comparison. The subsystem thermalization hypothesis works better for GGE than p-GGE, and better for typical states than eigenstates.
We finally consider the subsystem ETH [10, 11, 12], which has been extensively studied for integrable $(1+1)$-dimensional conformal field theories [17, 18, 19, 12, 20, 21, 22, 23, 24]. Here, we will consider the subsystem ETH for the non-integrable Ising chains. Since the conserved charge $\Pi$ commutes with the Hamiltonian $H$, thus the eigenstates are also the superselection states. As in the usual case for subsystem ETH, we first consider the demographics $S_{A}^{e}$ when comparing with the canonical ensemble thermal state, i.e., p-GGE specified by $\beta_{0}$, and the result is shown in Fig.9. We see the subsystem ETH holds well for $A=1$ and worsens as $A$ increases. We expect this $A$-dependence will be similar when considering the GGE or the other p-GGE. Thus, we now only consider the demographics of the subsystem ETH for only $A=1$. The demographics for GGE specified by $(\beta_{0},\beta_{1})$ and p-GGE by $\beta_{1}$ are shown in Fig.10, in which we also compare with their counterparts of typical-state thermalization hypothesis. Interestingly, for both eigenstates and typical states, the subsystem thermalization hypothesis works better for the GGE than the p-GGE specified by $\beta_{1}$. On the other hand, the subsystem thermalization hypothesis works better for the typical states than the eigenstates for a chosen type of thermal ensemble state. Our results verify the subsystem ETH in a broader context than the canonical ensemble and also show an interesting comparison with the typical-state thermalization hypothesis.

### v2

@@ II Thermalization of Quantum Systems with Conserved Charges
Figure 2: The evolving pattern of one-site entanglement entropy $S[\rho_{1}]$ of a $10$-site nonintegrable XXZ chain specified by (IV) and (22). This model has a $U(1)$ conserved charge denoted by $Q_{1}$, see (23). Each solid line represents the average of $S[\rho_{1}]$ over 500 states at each time step, which are evolved from the initial product states with fixed $\langle Q_{1}\rangle$. In this figure, we consider $\langle Q_{1}\rangle=0.0$ (red), $1.0$ (blue) and $2.0$ (green). The dashed lines are the values of $S[\rho_{1}]$ of the corresponding GGE states. The result shows that the values of $S[\rho_{1}]$ saturate to a non-thermal value around $T=25$.
@@
There is no unique definition of a typical state. In this work, we adopt the following definition used in [60]: its entanglement entropy of a single site with the rest of the spin chain, denoted by $S[\rho_{1}]$, is about the maximal value given by the time evolution of an initial product state. This will exclude some transient states, which may not be typical regarding entanglement spreading. We give an example for this construction by considering a ten-site nonintegrable XXZ spin chain with a $U(1)$ conserved charge $Q_{1}$. We then use ED to evolve a randomly generated product state with zero $S[\rho_{1}]$ but fixed $\langle Q_{1}\rangle$. In Fig. 2, we show the evolving pattern of the averaged $S[\rho_{1}]$ over $500$ initial states. The result shows that the (averaged) $S[\rho_{1}]$ values saturate to a maximum after $T\simeq 25$. However, these saturation values may differ from those given by the corresponding GGE states. Besides, the dependence on $\langle Q_{1}\rangle$ of these saturation values implies that the typical states can be classified by $\langle Q_{1}\rangle$. Each class corresponds to the almost $\{Q_{1}\}$-superselection sector discussed earlier. In this work, the typical states we collect for the demographics of the subsystem thermalization hypothesis are obtained by evolving randomly generated product states over a time duration of $T=1000$. This should be long enough to ensure the typicality. Based on this, we will prepare the typical states by choosing $T=1000$.
@@
III Thermalization demographics for quantum Ising spin chains with $Z_{2}$ symmetry
@@
As a starter, we consider the quantum Ising spin chain in the transverse field $h_{x}$ and longitudinal field $h_{z}$ with open boundary condition, given by the Hamiltonian
$H=\sum_{i=1}^{L}(h_{x}\sigma_{i}^{x}+h_{z}\sigma_{i}^{z})+\sum_{i=1}^{L-1}\sigma_{i}^{z}\sigma_{i+1}^{z},$ (17)
@@
where $\sigma_{i}^{x}$, $\sigma_{i}^{z}$ are Pauli matrix operator on the site $i$, and $L$ is the number of spins. This model possesses a $Z_{2}$ reflection symmetry $\Pi$, which swaps the order of spins by sending site $i$ to L+1iL+1-i, i.e.,
@@
$\Pi:=\begin{cases}P_{1,L}P_{2,L-1}\cdots P_{\frac{L}{2},\frac{L+2}{2}}&\quad\text{if }L=\text{ is even}\\ P_{1,L}P_{2,L-1}\cdots P_{\frac{L-1}{2},\frac{L+3}{2}}&\quad\text{if }L=\text{ is odd}\end{cases}\qquad$ (18)
where $P_{k,l}=(\sigma_{k}^{x}\sigma_{l}^{x}+\sigma_{k}^{y}\sigma_{l}^{y}+\sigma_{k}^{z}\sigma_{l}^{z}+I)/2$ is the permutation operator and $I$ is the identity operator. $P_{k,l}$ can swap the state of $k^{th}$ and $l^{th}$ sites. To avoid introducing additional conserved charge due to translational invariance, in this work, we will consider open spin chains without imposing periodic boundary conditions.
@@
Figure 3: Patterns of typical-state thermalization demographics of $S^{t}_{A=1}$ of a $10$-site Ising chain of (17) near an integrable point when comparing with the corresponding GGE state specified by the chemical potentials $(\beta_{0},\beta_{1})$ of the conserved charges (20). The phase diagram of the Ising chain is shown in the inset of the bottom subfigure, where a red line indicates the region we plot for thermalization demographics. The integrable point on the red line is $h_{x}=1.05$ and $h_{z}=0$. We plot $10$ demographics from $h_{z}=0$ to $h_{z}=0.1$. At the integrable point, the GGE state is simply a p-GGE state, as all conserved charges other than $Q_{1}$ are neglected. The results show that the thermalization hypothesis breaks down significantly near the integrable point, but gradually holds better as one moves away from it.
@@
This model is integrable if $h_{x}=0$ or $h_{z}=0$. Otherwise, it is nonintegrable. Its phase diagram is shown in the inset of the bottom subfigure of Fig. 3, e.g., see [64]. It was shown that the level-spacing statistics of the superselection sectors with $\Pi=\pm 1$ match with the Wigner-Dyson distribution for the nonintegrable Ising chain [60]. Despite that, we can examine how the pattern of the demographics of $S^{t}_{A}$ changes by tuning the coupling constant to make the model more nonintegrable. The thermalization hypothesis holds well if most of the population lives near $S_{A}^{t}=0$. Otherwise, we can quantify the violation of the thermalization hypothesis by the spreading of the population away from $S_{A}^{t}=0$. A typical result for such pattern changes is shown in Fig. 3. We observe that the demographics of $S_{1}^{t}$ for the integrable case significantly violate the subsystem thermalization hypothesis. However, as the model moves away from the integrable point, the values of $S_{1}^{t}$ gradually converge to zero.
From now on, we will only consider the thermalization for the non-integrable variants of the quantum spin chains. Thus, for the Ising chain, we will set the parameters to the following:
@@
We will present the demographics of $S_{A}^{t,a,e}$ for various types of p-GGE states.
@@ III.1 Demographics of typical-state thermalization
Figure 5: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{0}$ for the Ising chain of (19) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, as seen for $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{1}\rangle$.
@@
We first consider the subsystem thermalization hypothesis for the typical states. In Fig. 4, we show the thermalization demographics of the relative entropy $S^{t}_{A}$ of the region $A$ with $1$, 33, 55 sites (denoted by $A=1,3,5$) between a typical state and the corresponding GGE specified by $(\beta_{0},\beta_{1})$. Since the GGE contains two chemical potential parameters $\beta_{0}$ and $\beta_{1}$, we present the demographics in the form of density plots. For $A=1$, $S^{t}_{A}$’s of all the typical states considered are almost vanishing. This implies that the thermalization hypothesis for typical states is effective for $A=1$ when compared with the corresponding GGE thermal states. On the other hand, for $A=3$, the thermalization hypothesis still looks good, with some exceptions for large $|\beta_{0}|$ or $|\beta_{1}|$. However, when $A=5$, the thermalization hypothesis breaks down. The above patterns of dependence of validity of the thermalization hypothesis on the size $A$ are expected and also give an explicitly quantitative picture. Moreover, in Appendix B, we also performed the size-scaling study by varying $L$ to show that the thermalization hypothesis works better for larger systems. This study is not quite extensive due to the limitation of computing power, but the tendency, however, is clear in meeting the expectation.
@@
Besides checking the typical-state thermalization hypothesis, Fig. 4 also reveals the demographics of the typical states on the $(\beta_{0},\beta_{1})$ of the corresponding GGE ensemble states. Interestingly, the inverse temperature $\beta_{0}$ can be either positive or negative, and it is almost symmetric about $\beta_{0}=0$. The GGE states with negative (inverse) temperatures are hypothetical and cannot be realized naturally. However, we see that they can be realized by typical states. Moreover, the thermalization hypothesis works even for such hypothetical GGE thermal ensemble states. On the other hand, the chemical potential $\beta_{1}$ for reflection symmetry $\Pi$ is almost negative for the typical states we construct. It is unclear why there is such a bias.
@@
After considering the comparison with GGE states, we now compare with the corresponding p-GGE thermal states. In Fig.5, we present the typical-state demographics of $S^{t}_{A}$ when compared with the corresponding p-GGE states specified only by $\beta_{0}$ (inverse temperature). On the other hand, in Fig. 6, we compare the typical states with the corresponding p-GGE states specified only by $\beta_{1}$. We see that the features of the demographics of p-GGE on either $\beta_{0}$, i.e., symmetric about $\beta_{0}$, and $\beta_{1}$, i.e., favors negative values, are the same as the ones of GGE. Regarding the validity of the typical-state thermalization hypothesis, we see in both Fig. 5 and 6 the similar patterns of dependence on the size $A$ as in Fig. 4 for GGE states’ comparison. In all three cases, the thermalization hypothesis for the typical states works quite well for $A=1$ and gradually worsens with increasing $A$.
@@
Although it is generally believed that the typical-state thermalization hypothesis will hold for small subsystems, we emphasize that it is unclear whether it will hold for various thermal ensemble states and how local the subsystem must be for it to be applicable. The results presented here clarify these issues.
@@ III.2 Demographics of almost superselection-state thermalization
In Fig. 7, we consider the almost superselection typical states with fixed $\langle Q_{1}\rangle$ and compare them with the corresponding p-GGE states specified by $\beta_{0}$. In Fig. 8, we swap the role of the $Q_{0}$ and $Q_{1}$, that is, compare the almost superselection states with fixed $\langle Q_{0}\rangle$ to the corresponding p-GGE states specified by $\beta_{1}$. In both cases, we show the population ratio demographics against $S_{A}^{a}$. The population ratio for a given $\langle Q_{a}\rangle$ is defined with respect to the total number of all typical states considered. We can quantify the violation of the thermalization hypothesis by the spreading of the population away from $S_{A}^{a}=0$. In the first case, as shown in Fig. 7, the thermalization hypothesis holds well for each almost superselection sector if $A=1$. Moreover, the patterns of the populations for different $A$’s do not change much as $\langle Q_{1}\rangle$ varies. In the second case, as shown in Fig. 8, the thermalization holds well only for $\langle Q_{0}\rangle=0$ and worsens more and more as $\langle Q_{0}\rangle=0$ increases, even for $A=1$. In this case, even the typical-state thermalization holds, but its fine-grained version does not. Unlike in Fig. 7, the patterns of populations for different $A$’s change more significantly as $\langle Q_{0}\rangle$ varies.
It may look strange why the coarse-grained version of the thermalization hypothesis holds, but most of its fine-grained versions fail. However, if we look into the scales of the population ratio in each subfigure, it can be understood that the population respecting the thermalization hypothesis dominates; see also Fig. 6(b), where the sector of $\langle Q_{0}\rangle=0$, which obeys the thermalization hypothesis as shown in the first subfigure of Fig. 8, has the largest population.
@@ III.3 Demographics of eigenstate thermalization
Figure 9: Subsystem eigenstate thermalization demographics: $S^{e}_{A=1,3,5}$ vs. $\beta_{0}$ for the Ising chain of (19) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, i.e., $A=3$ (blue circles) and $A=5$ (black triangles).
Figure 10: Subsystem eigenstate thermalization demographics of $S^{e}_{A=1}$ for the Ising chain of (19) by comparing with (i) GGE states (empty red) and (ii) p-GGE states specified by $\beta_{1}$ (solid pink). We also present the typical-state counterparts (empty blue and solid gray) for comparison. The subsystem thermalization hypothesis is more effective for GGE than p-GGE, and more applicable to typical states than to eigenstates.
We finally consider the subsystem ETH [11, 12, 13], which has been extensively studied for integrable $(1+1)$-dimensional conformal field theories [22, 23, 24, 13, 25, 26, 27, 40, 41]. Here, we will consider the subsystem ETH for the non-integrable Ising chains. Since the conserved charge $\Pi$ commutes with the Hamiltonian $H$, the eigenstates are also the superselection states. As in the usual case for subsystem ETH, we first consider the demographics $S_{A}^{e}$ when comparing with the canonical ensemble thermal state, i.e., p-GGE specified by $\beta_{0}$, and the result is shown in Fig.9. We observe that the subsystem ETH performs well for $A=1$ and deteriorates as $A$ increases. We expect this $A$-dependence will be similar when considering the GGE or the other p-GGE. Thus, we now only consider the demographics of the subsystem ETH for $A=1$ only. The demographics for GGE specified by $(\beta_{0},\beta_{1})$ and p-GGE by $\beta_{1}$ are shown in Fig.10, in which we also compare with their counterparts of the typical-state thermalization hypothesis. Interestingly, for both eigenstates and typical states, the subsystem thermalization hypothesis works better for the GGE than the p-GGE specified by $\beta_{1}$. On the other hand, the subsystem thermalization hypothesis works better for typical states than for eigenstates of a chosen type of thermal ensemble state. Our results verify the subsystem ETH in a broader context than the canonical ensemble and also show an interesting comparison with the typical-state thermalization hypothesis.